Y Company obtained the following balances from its computerized accounting information system at the end of the year before​ adjustments:

Accounts receivable ​$ ​ 27,000

Allowance for uncollectible accounts ​ (3,000)

Net sales   ​100,000

Bad debts expense 0

The company estimates that​ 20% of accounts receivable will be uncollectible. After the correct adjusting entry has been​ made, the allowance for uncollectible accounts will​ be:

Y Company obtained the following balances from its computerized accounting information system at the end of the year before​ adjustments:

Accounts receivable                                                                                                                      ​$ ​ 27,000

Allowance for uncollectible accounts ​ (3,000)

Net sales                                                                                                                                                                                   

  ​100,000

Bad debts expense                                                                                                                                                                          0

The company estimates that​ 20% of accounts receivable will be uncollectible. After the correct adjusting entry has been​ made, the allowance for uncollectible accounts will​ be:

A.

​$(4,800).

B.

​$(2,400).

C.

​$(20,000).

D.

​$(5,400).

The company estimates that​ 2% of net sales will be uncollectible. After the correct adjusting entry has been​ made, what is the Net accounts receivable on the

year−end balance​ sheet?

A.

​$22,000

B.

This is a trick question. Accounts receivable appear on the income​ statement, not the balance sheet.

C.

​$24,000

D.

​$27,000
Answer #1

1) Allowance for doubtful account balance after adjusting entry = Account receivable*percent uncollectible

= 27000*20%

Allowance for doubtful account balance after adjusting entry = 5400

So answer is d) ($5400)

2) Bad debt expense = 100000*2% = 2000

Net account receivable = 27000-5000 = 22000

So answer is a) $22000